In the reaction below, 8.0 g of H2 react with 9.0 g of O2. Which of the following statements is true?
_2_ H2 + _1_ O2 _2_ H2O
1
Expert's answer
2020-04-15T01:52:31-0400

Here, the reaction is as follows

"H_2+O_2\longrightarrow2H_2O"

Here given that mass of H2=8.0 g and mass of H2O= 9.0 g

moles of "H_2=\frac{given mass}{molar mass}= \frac{8}{2}= 4" mole and


moles of "O_2= \frac{given mass}{molar masS}=\frac{9}{32}" =0.2812


so here O2 is limiting reagent and

massof "H_2O = 18\times.2812=5.06 g"